Suppose the economy is operating at equilibrium, with Yo=1,000. If the government undertakes a fiscal change whereby the tax rate, t , increases by .05 and government spending increases by 50, will the budget surplus go up or down? Why?
1
Expert's answer
2021-08-13T13:57:52-0400

Budget surplus = Tax revenues - Government expenditure

Given that economy is operating at equilibrium where Yo = 1000

Since the tar rate is 0.05 or 5%

The amount of change in tax "5\% of 1000 = \frac{5}{100}×1000 = 50"

The amount of increase in the government expenditure is 50

Budget surplus = 50 - 50 = 0

Based on the reasoning and the above calculation we can conclude that budget surplus will neither go up nor go down as it will remain constant because both the government expenditure and government revenue are increasing by the same amount